When performing a hypothesis test, the p value is found to be 0.031. This means that:

a) Ho is rejected at the level α = 0.05 but not at the level α = 0.01

b) Ho is rejected at the level α = 0.05, and at the level α = 0.01

c) Ho is not rejected at levels α = 0.05 and α = 0.01
Answer #1

Solution,

P-value = 0.031

correct option is = a

\alpha = 0.05  

P-value < \alpha

Reject the null hypothesis .

and

\alpha = 0.01

P-value > \alpha

Fail to reject the null hypothesis .